OpenStudy (anonymous):

Stephanie receives a salary of $650 per month plus a commission of 5.5% on the first $3,000 of sales, and 7% of all sales over $3,000. Find Stephanie's gross pay for the month if her sales were $10,890.

OpenStudy (anonymous):

is this right? 650+(3000*.055)+((11230-3000)*.7) 650+165+(8230*.7) 650+165+576.1 1391.10

OpenStudy (anonymous):

Where did you get 11230 from? Here's my working: 650 + 165 + ((10890-3000)/100*7) = 650 + 165 + (7890/100*7) = 650 + 165 + (552.3) = 1367.3 So her total pay is $1367.30

OpenStudy (anonymous):

I'm taking an online study exam and didn't realize it changed the numbers when i refreshed the page. I understand your process now, thank you!